# S-EQA: Tackling Situational Queries in Embodied Question Answering

### v1

@@ III-C1 Analyzing S-EQA with t-SNE
III-C2 Assessing S-EQA’s Utility with MTurk
We conduct a large-scale Mechanical Turk [7] study to gauge S-EQA’s authenticity and usability in answering real-world queries.
In our first experiment, we assess whether the consensus object data being generated adequately facilitates answering situational queries. Workers are shown a query, images of the corresponding room, and images of objects set to consensus states, while the LLM Answer remains hidden to negate bias. Given options of Yes, No, and Cannot Answer, a worker’s capacity to give a definitive response, i.e., choosing Yes or No, serves as an indicator of whether the generated data is in alignment with general user consensus. For instance, the query “Is the house ready for sleeptime?”, if unanimously agreed upon by workers when shown images of dimly-lit rooms, closed doors, and windows, validates the data point.
Subsequently, in a second experiment, workers are presented with an image of a room accompanied by a situational S-EQA query (For eg. Is someone working in the bedroom? with top-views of the bedroom) and, separately, images of consensus objects paired with simpler, objective queries about their state (For eg., Is the Computer On? with an image of a computer being on). For the simple query annotation, we additionally sample ‘Yes/No’ questions the SimVQA [24] dataset to even out the number of situational and simple queries. An improved annotator performance in simple over situational queries would indicate S-EQA’s real-world usability in providing human-authenticated consensus object-states for situational queries which are relatively simpler to answer.
III-D Evaluating S-EQA in VirtualHome
Assuming low-level navigation to be solved, we quantitatively evaluate S-EQA via Visual Question Answering (VQA). VirtualHome provides us with a household environment where object states visually change upon modifying states. We present VQA models with the same setup as described in our first MTurk experiment presented in III-C2, and compare the results. Figure 4 presents an overview.
Room VQA (Situational Query): Here, we present the VQA model with multiple top-view images of a relevant room as input, and ask it to answer the associated situational query with Yes/No/Cannot Answer. Most of our queries are associated with one particular room (see Figure 3(a)), which is often part of the query itself. During our experimentation, we observed a clear bias in model performance with such queries. For instance, if the query contained ‘living room’, and the image was that of the living room, the answer would always be yes, irrespective of the actual context of the query. To mitigate this, we replace any mention of a room in the query with the generic term - “this place”.
Object VQA (Simple Query): Here, we use consensus object states linked with the situational query to formulate simple objective questions for a VQA model. These are of the form -
@@ III-D Evaluating S-EQA in VirtualHome
We provide the VQA model with questions along with images of the consensus objects in their designated states.
@@
and give us an upper bound on the performance of VQA models on S-EQA in VirtualHome. For , we set a threshold of above which the score would be deemed satisfactory.
IV Results and Inferences
We use GPT-4 [12] as the LLM of choice to generate situational queries in our experiments. Our ablations with GPT-3.5 showed poor quality in terms of generation, with many outputs not fitting into our specified datapoint format.
IV-A Preparing the Simulator
We first program the object states and relationships from the generated data points into the VirtualHome environment. This also includes filtering out a small subset of bad states which are not feasible. For instance, ‘computer INSIDE fridge’ is not a valid relationship for us to set, and the simulator prevents this.
For VQA and the MTurk user study, we require images of the room and corresponding consensus objects set to their appropriate states. We gather room images by considering the 4 corners of each room, with a camera pointing downwards towards the center. Choosing images in this way allows us to achieve maximum coverage of objects present.
Then for object images, we use the first-person view of a human placed in the environment. The human is instructed to walk to each of the objects present in a S-EQA datapoint, and we grab an image when they are present in front of them. This approach for capturing images proved simpler to implement and is in-line with the visual perspective of an agent grabbing an object.
IV-B Quantitative Results and Benchmark
We use BLIP-2 [25], BLIP [26] and ViLT [27], all finetuned on VQAv2 [28] to evaluate S-EQA, and these results are presented in Table II.
@@ IV-B Quantitative Results and Benchmark
Observe the higher Object-VQA accuracies and F1 scores over Room-VQA across each approach. This suggests that answering situational queries by looking at images at a room level is significantly harder than answering simple queries about the states of target objects. Our second MTurk experiment (Refer III-C2) further corroborates this result, suggesting S-EQA’s usability in providing the simpler consensus object data for answering situational queries.
The high values on joint scores indicate an upper bound on performance. While still better than RoomVQA, ObjectVQA models have mediocre performance overall, despite the queries being simple. This can be attributed to inference on synthetic images from VirtualHome, which these models have not been trained on. Our real-world experiments suggest improved performance, confirming this domain gap.
@@
IV-C Qualitative Results and Inference
In both our MTurk experiments (Refer III-C2), we measure average reviewer agreement across 5 users on each question to obtain ground truth labels on the data. For instance, if a question has Yes’s, and No’s, we label the ground truth as ‘Yes’. On the other hand, if even one of the annotations is ‘Cannot Answer’, we label the human-annotated output by the same.
For the first experiment, we obtain a low ‘Cannot Answer’ percentage of among the nearly participants. As such, a high percentage () of generated queries being answerable strongly indicates that the generated consensus object states are good representation of the human consensus for each query. Conversely, a low correlation of in LLM-generated and worker-annotated answers suggests the LLM’s limited capability in adequately answering them. To analyze this further, we attempt to reason about the generated answers by prompting GPT-4 to provide reasoning:
@@ IV-C Qualitative Results and Inference
The reasoning generated for the LLM Answer is as follows -
@@
In our second MTurk survey, we sought to assess the usability of S-EQA by comparing situational and simple queries. For simple queries, of participants answer accurately, indicating their mostly unambiguous and straightforward nature. In contrast, S-EQA queries show only a accuracy. of users also choose Cannot Answer on the situational queries, as opposed to on the simple ones. This discrepancy showcases the usability of S-EQA in providing consensus objects and states for indirectly answering situational queries.
We infer from these qualitative experiments that while GPT-4 is indeed capable of producing situational queries and authentic consensus information, it often goes against the norm in answering the query itself. As such, an agent equipped with an LLM for directly answering situational queries from scene images might not be able to do so correctly. S-EQA is very useful here, as it simplifies the QA task by providing authentic objects and states relevant to the situational query for the agent to ground and subsequently answer the query.
V Conclusion, Limitations and Future Work
We tackle the novel and challenging problem of Embodied Question Answering with Situational Queries. Unlike prior work that dealt with simple objective queries pertaining target objects, situational queries necessitate an approximation of a consensus on object states. We use a Large Language Model (LLM), GPT-4 to approximate this consensus, incorporating it in an approach to pose unique situational queries, corresponding consensus object information and predicted answers. We use VirtualHome as an embodied simulation platform for our experiments, as it provides us with modifiable object states that also visually change in their appearance upon modification, allowing for quantitative validation.
We validate the authenticity of the generated data via a large scale MTurk user study with participants, and introduce it as S-EQA, the first EQA dataset addressing situational queries. In the study, we observe a low correlation of the GPT-4’s answers with that of the user, and notice a very high percentage of users () believe the queries are answerable with the provided consensus states. This lets us infer that while GPT-4 is capable of reliably generating consensus objects and states required for answering a situational query, it is incapable of directly answering the situational query itself. From this, we infer S-EQA’s usability, in providing a human-verified dataset of consensus data that could be used as a prior to break down situational queries into simpler, objective ones based on a human-consensus to answer them in household environments. Further, we perform Room and Object level VQA on S-EQA and quantitatively observe this discrepancy, and setting a benchmark for validating S-EQA on VirtualHome.
As a future experiment, we will look at a real-world deployment of an S-EQA-powered household robot, to gather data and finetune the LLM’s generational capabilities. The queries in S-EQA are limited due to the abstract, situational nature of the questions being generated and the limited number of objects. Future work will explore cases when the user poses novel questions outside the S-EQA dataset, requiring semantic alignment of the novel query with those in the dataset to extract consensus objects and states for S-EQA.

### v3

@@ III-C Validating S-EQA: t-SNE plots and M-Turk Annotation
MTurk Validation: We conduct a large-scale Mechanical Turk [8] study to annotate the generated data and evaluate the authenticity of real-world queries. The annotators are chosen from the US-East region, with at least a approval rate on previous work.
Annotators are shown a generated query, along with consensus object-states and relationships. Given options of Yes, No, and Cannot Answer, a worker’s capacity to give a definitive response, i.e., choosing Yes or No, serves as an indicator of whether the generated data is in alignment with general user consensus. For instance, the query “Is the house ready for sleeptime?”, if agreed upon by workers when shown images of dimly-lit rooms, closed doors, and windows, validates the data point.
We obtain a low ‘Cannot Answer’ percentage of among the participants. As such, a high percentage of generated queries being answerable strongly indicates that the LLM is capable of generating object-states and queries that are a good representation of human consensus.
Subsequently, we conduct a second comparison study with the same workers, where they are presented with an image of a room accompanied by a situational S-EQA query (For eg. Is someone working in the bedroom? with top-views of the bedroom) and, separately, images of objects paired with consensus queries formulated based on their state (For eg., Is the Computer On? with an image of a computer being on). We sample additional ‘Yes/No’ questions the SimVQA [24] dataset to even out the number of situational and consensus queries. An improved annotator performance in answering consensus queries over situational ones would indicate S-EQA’s utility in breaking down subjective queries into simpler consensus queries for improved question-answering. We discuss this later in the results section.
In both our MTurk experiments, we measure average reviewer agreement across 5 users on each question to obtain ground truth labels on the data. For instance, if a question has Yes’s, and No’s, we label the ground truth as ‘Yes’. On the other hand, if even one of the annotations is ‘Cannot Answer’, we label the human-annotated output by the same.
IV Situational EQA: Evaluation
Figure 5 presents an overview of our evaluation of the simulator-based dataset, and Figure 6 presents our real-world setup for evaluation. We evaluate queries both via VQA as well as by using an LLM. Further, LLM answers are qualitatively studied via reasoning to infer situational awareness.
IV-A VQA Evaluation:
We perform VQA on images taken from VirtualHome with situational and consensus queries. The VirtualHome simulator is unique in enabling this, as it allows us to modify consensus object states pertaining to subjective queries such that they also visually change upon modifying states.
Situational Consensus Queries: Consensus object states associated with the generated situational queries are used to formulate binary objective questions for evaluation. These are of the form -
@@ IV-A VQA Evaluation:
For instance, for a query ‘Was someone working in the bedroom?’, the consensus queries would be ‘Is the Computer On?’ and ‘Is the Lightswitch On?’.
We present VQA models with both situational and consensus queries, and compare the results.
Room VQA (Situational Query): Here, we present the VQA model with multiple top-view images of a relevant room as input, and ask it to answer the associated situational query with Yes/No/Cannot Answer. The assumption here is that the agent might have access to overhead security cameras which it could use to answer the queries. Most of our queries are associated with one particular room, which is often part of the query itself. During our experimentation, we observe a bias in model performance when the name of the room itself is present in the query. For instance, if the query contained ‘living room’, and the image was that of the living room, the answer would always be yes, irrespective of the actual context of the query. To mitigate this, we replace any mention of a room in the query with the generic term - “this place”.
Object VQA (Consensus Query): We provide the VQA model with these consensus queries along with images of the corresponding objects in their designated states.
@@
and give us an upper bound on the performance of VQA models. For , we set a threshold of above which the score would be deemed satisfactory.
IV-B LLM Evaluation:
Motivated by K-EQA [7], we utilize scene graphs provided by VirtualHome to develop an LLM-based action planner.
For S-EQA, we have situational queries generated along with consensus object states and relationships. VirtualHome provides us with a fully observable scene graph for this, i.e., the states and relationships of all objects are known.
Our aim is to measure how well the LLM is able to approximate the human consensus behind answering situational queries. For this, we first ask the LLM to provide a Yes/No answer to each of the the datapoints generated. We then measure the correlation between the LLM and human annotated answer. Further, we also ask the LLM to reason about its answers for a subjective analysis.
IV-C Real World Experiments:
Unlike VirtualHome, where a fully observable scene graph is readily available, the real-world environment does not have object states in a predefined format. Instead, we gather images from various locations in our environment which is used as context for GPT-4V to generate situational datapoints. A representative subset of scenes and datapoints being generated are illustrated in Fig 6. We aimed to create diverse situations including group meetings, robot operations, cafeteria related events across various locations in our lab environment.
The VLM predicts Yes/No responses for each datapoint. These are then compared against human-annotated ground truth responses to assess alignment. We then analyze failure cases where the LLM misinterprets missing information or hallucinates object states, showcasing the difficulty in accurately comprehending situational queries in the real world.
V Results and Inferences
GPT-4 [13] is our LLM of choice to both generate and evaluate situational queries in our experiments. Our ablations with GPT-3.5 showed poor quality in terms of generation, with many outputs not fitting the specified data format.
V-A Simulator Setup
We first program the object states and relationships from the generated data points into the VirtualHome environment. This also includes manually filtering out a small subset of bad states which are not feasible. For instance, ‘computer INSIDE fridge’ is not a valid relationship for us to set, and the simulator prevents this.
For VQA evaluation, we require images of the room and corresponding consensus objects set to their appropriate states. We gather room images by considering the 4 corners of each room, with a camera pointing downwards towards the center. Choosing images in this way allows us to achieve maximum coverage of objects present.
Then for object images, we use the first-person view of a human placed in the environment. We consider the human to be our robot agent here. They are instructed to walk to each of the objects present in a S-EQA datapoint, where we grab an image when they are present in front of them. This approach for capturing images proved to be simpler to implement over simulator camera placement and gives us a visual perspective of an agent grabbing an object.
V-B VQA Results
@@
Observe the higher Object-VQA accuracies ( on average) and F1 scores over Room-VQA across each approach. This suggests that answering situational queries by looking at images at a room level is significantly harder than answering queries about the states of target objects. Our comparison study with MTurk (Refer III-C) further corroborates this result, with of users finding consensus queries to be simpler to answer over situational ones. This suggests that our approach of breaking down situational queries into simpler consensus ones could be used to indirectly answer difficult situational queries.
The high values on joint scores indicate an upper bound on performance. While still better than RoomVQA, ObjectVQA models have mediocre performance overall, despite the queries being straightforward in nature. This can be attributed to inference on synthetic images from VirtualHome which these models have not been trained on, opening an avenue for future work.
Language - Question length, re-classification percentage, Plot curve of average length of dataset questions. VQA - Room VQA and Object VQA
@@
V-C LLM Eval Results
We obtain a low correlation of between GPT-4 generated answers and the human-annotated ones. This is despite the LLM having complete access to the entire scene graph which it used earlier to generate situational queries.
Furthermore, the low ‘Cannot Answer’ percentage of on the annotated responses shows that most queries and consensus data generated did indeed make sense to most users. We can infer from these two values that while GPT-4 is good at generating situational queries and the potential consensus data needed to answer them, it is poor at answering it’s own generated queries. To understand this further, we prompt for reasoning as follows:-
LLM Reasoning: To analyze the poor LLM performance on query answering, we reason about the generated answers by prompting GPT-4 to provide reasoning as follows:
@@ V-C LLM Eval Results
The reasoning generated for the LLM Answer is -
@@
These results suggest that while GPT-4 is indeed capable of producing situational queries and authentic consensus information, its displays poor commonsense reasoning while answering such queries.
V-D Real World Experiments: Results and Discussion
We experiment with situational queries in our real-world lab environment, assuming the presence of an assistive workplace robot agent. We gather a total of images across various situations, corresponding to different locations in our workplace, and feed these into PGE as visual context for GPT-4V. We then generate datapoints using PGE and manually annotate them.
During annotation, we note several hallucinations in the object states that did not occur earlier when the scene graph was explicitly defined in VirtualHome. Some instances of this are highlighted in red in Figure 6. For example, note the noise level indication is present as an object state in many cases, but can not be discerned by looking at the image. Ghost objects such as microphones and safety equipment also appear in the generated object-states. We infer here that a grounded scene graph, like what VirtualHome provides is important to avoid poor data generation.
Evaluating the generated datapoints with GPT-4 presents a mismatch of , a similar trend as that with the VirtualHome data, and solidifies our claim that LLM generated binary answers are misaligned with human consensus.
Similar to what was done with VirtualHome, we conduct LLM based reasoning on the queries generated. An example of this is illustrated in Fig. 8. Note the counterintuitive explanation that GPT gives when asked if the meeting space is ready for a group discussion. Despite the table being occupied with someone present in the room, who also appears to be using their laptop (Object State includes Laptop:OPEN), the reasoning is completely off. We observe many such cases in the data that we gather, where LLM answers and the justification provided goes against human consensus.
LLMs have been widely adopted in robotics for their strong commonsense reasoning, enabling significant performance gains across embodied tasks [18, 16, 28]. These models have contributed to the rise of foundation models for robotics, where a common paradigm involves LLMs predicting task outputs or structured action plans, which agents then execute with varying affordances [29]. Given these successes, one would expect LLMs to exhibit similar strengths in our task.
However, we observe a counterintuitive outcome—LLMs perform poorly in commonsense reasoning. While this may stem from the inherent subjectivity of situational queries, it also reveals GPT-4V’s limited situational awareness, even when provided with visual cues. Additionally, hallucinations in object states likely degrade downstream query performance, highlighting how even minor inaccuracies in state identification can significantly impact task execution. This contrasts with prior work where LLMs have demonstrated strong object reasoning, suggesting that situational awareness presents a distinct challenge. Future work will look at improving LLM-based situational reasoning by employing more robust multimodal alignment schemes via image captioning, memory-aware mechanisms, or explicit uncertainty modeling to mitigate state hallucinations and enhance task reliability.
VI Conclusion, Limitations and Future Work
We introduce and tackle the novel and challenging problem of Embodied Question Answering with Situational Queries — S-EQA. Unlike prior work that dealt with simple objective queries that directly reference target objects, situational queries necessitate an approximation of a consensus on multiple object states. Given the challenge in gathering such queries from the real world, we rely on commonsense knowledge from an LLM, GPT-4 to develop a Prompt-Generate-Evaluate (PGE) framework for posing unique situational queries and corresponding consensus object information. In the VirtualHome simulator, we use PGE to generate 2K unique queries, and annotate them via a large scale MTurk user study with participants. The high answerability () lets us infer that LLMs are good at generating situational queries and corresponding consensus data.
However, using GPT-4 to answer the simulator queries shows a low correlation with the human annotations (); we infer that LLMs are good at generating situational data but poor at answering them. When asked to explain its answers, we observe the LLM’s poor commonsense capability, exhibiting counterintuitive reasoning. This suggests that LLMs show poor situational awareness.
Finally, we discuss challenges with transferring our generation and evaluation schemes to a real world lab environment. Since a grounded scene graph isn’t directly available for use in the real-world, we provide images of various locations as context for query generation, and observe several unintended hallucinations with ghost object states as consensus. Future work will focus on limiting these by first gathering structured scene graphs using existing visual map building techniques.